OpenStudy (anonymous):

2cosx=2sinxcosx Why?

OpenStudy (anonymous):

it is possible when angle become 90.

OpenStudy (anonymous):

Ok, new question. differentiate this with respect to x \[\sin^2x\]

OpenStudy (anonymous):

It involves this question

OpenStudy (anonymous):

2sinxcosx.. ^_^

OpenStudy (dumbcow):

it equals 2sincos

OpenStudy (anonymous):

Ok. How?

OpenStudy (anonymous):

by rule of dervative

OpenStudy (dumbcow):

chain rule u = sin x du = cos x f(u) = u^2 f'(u) = 2u --> 2u*du = 2sincos

OpenStudy (anonymous):

yeah.. rock . roll

OpenStudy (anonymous):

Oh Ok. I thought you just switched the sin to cosin. But I guess, since it's squared... that rule doesn't apply fully.

OpenStudy (anonymous):

u = (sinx)^n f(u) = n(u)^(n-1)du